A real estate agent drove to a remote property. She averaged 50 mph to the property and 48 mph on her return trip. Her total driving time was 4 hours and 54 minutes. How far was the property?
Answer by mananth(12270)

x/50+x/48 = 4 9/10 hours => 49/10 hours
LCD =50*48.multiply by the LCD
(48x+50x)/2400=49/10
98x/2400=49/10
cross multiply
98x=2400*49/10
98x=2400*49/10
x= 2400*49/98*10
x= 120 miles